### v2

@@ IX Radiometeorology Mixed URSI, IUGG, IUPAP
2727footnotetext: Data taken from “Report 5th General Assembly, September 1947," p. 5, Series B2aa, Vol. 1, IuG. In 1947 the explicit mention of type General had not yet been introduced, and the types Restricted and Grand was later renamed Specialized, and Affiliated.
In the wake of the 1947 General Assembly, the proliferation of commissions marked a significant phase of growth for IUPAP, but the scope and function of these commissions initially lacked clarity and differed from their eventual roles. Prior to 1947, IUPAP’s structure mainly consisted of general-purpose commissions, such as the SUN Commission and the Commission on Publications. The 1947 General Assembly saw the establishment of several new commissions alongside the reconstitution of the SUN Commission, which was inactive since the mid-1930s (see Table 1). These newly established commissions fell into three categories. The first, known as “mixed commissions," were inter-union commissions created by ICSU to tackle urgent issues by linking multiple scientific unions, including IUPAP, which ratified their creation and the participation of its members. The second category, termed “grand commissions," focused on distinct sub-fields of physics. The first of this type was the International Commission for Optics (ICO), founded during the 1947 General Assembly. [^28] For the foundation of the ICO, see [24]. Its secretary was Pierre Fleury, a prominent French optical physicist who was also appointed General Secretary of IUPAP at the same meeting. The establishment of the ICO set the precedent for the structure of grand commissions: they were independent international scientific entities that became integrated within IUPAP to pursue their objectives under its auspices. Grand commissions would immediately be renamed “affiliated commissions", which exist to this day, but had a limited development at the time, given the fact the the second affiliated commission was established only in 1974.292929For the establishment of the second IUPAP affiliated commission, see [31]. The third category, “restricted commissions," included the Commission on Thermochemistry (commission III), the Commission on Cosmic Rays, and the Commission on Radioactive Units. These commissions were envisioned to pursue a relatively “limited program" compared to the broader scope of mixed and, especially, grand commissions, which aimed to forge specific physics sub-fields in the international arena. Some restricted commissions were anticipated to evolve into either mixed or grand commissions. For instance, the Commission on Radioactive Units soon transitioned into the Joint Commission on Radioactivity, a collaborative effort between IUPAP and IUPAC. [^30] For an historical analysis of the Joint Commission on Radioactivity, see [17].
@@
In January 1948, under the auspices of IUPAP a first scientific meeting was organized by Prigogine at the Free University of Brussels. Entitled Symposium on Thermodynamics, the Brussels meeting saw the participation of 22 Western European chemists and physicists. They convened to explore topics in statistical thermodynamics, cryogenics, and irreversible processes.333333“IUPAP, Minutes Sixth General Assembly (1948),” p. 6, Series B2aa, Vol. 1, IuG. The deliberations and outcomes of this meeting were later disseminated in the conference proceedings, thanks to the support of UNESCO [48]. This event marked a foundational step in the Commission’s expanded role and scope, evolving from a body primarily focused on supporting the SUN Commission’s standardization efforts in thermodynamics to a thematic commission. This new role designated it as an IUPAP commission dedicated to fostering the development of a physics sub-discipline on an international scale, primarily through the cultivation of a dedicated scientific community. The Brussels conference, as the Commission’s inaugural activity, underscored this strategic pivot. The chosen approach to achieve international community-building was the orchestration of conferences centered on specialized research areas.
@@
This modification of the Commission’s functions was ratified at the sixth IUPAP General Assembly, convened in Amsterdam in July 1948, marking the second assembly after the end of World War II. At this meeting, the composition of the Commission was significantly broadened. Alongside Edward A. Guggenheim, who served as president, and Ilya Prigogine, who took on the role of secretary, the original members, Bauer and de Boer, were joined by two scientists from the United States: physical chemist James Alexander Beattie from MIT and chemical physicist Joseph Edward Mayer from the University of Chicago.343434“IUPAP, Minutes Sixth General Assembly (1948),” p. 6, Series B2aa, Vol. 1, IuG. This inclusion of U.S. scientists represented a notable shift, as they had been absent from the first post-World War II IUPAP General Assembly and so became active in IUPAP’s commissions only from the Amsterdam General Assembly onward.
@@
The decisions regarding the Commission’s revamped function proved to be even more pivotal. Following Prigogine’s presentation on the Commission’s activities, the General Assembly endorsed a significant expansion of its responsibilities and introduced a new designation: “Commission on Thermodynamics and Statistical Mechanics."353535“IUPAP, Minutes Sixth General Assembly (1948),” p. 13, Series B2aa, Vol. 1, IuG. The narrower focus on defining magnitudes and notations within thermodynamics was agreed to continue under the collaboration with the SUN Commission. This effectively reinstated the SUN Commission’s role in standardizing thermodynamic measures, as evidenced by its report at the 1948 General Assembly. [^36] Ibid., p. 6. Furthermore, there was an ambitious plan to host an international symposium addressing current challenges in statistical mechanics. This event was tentatively scheduled for 1949, with the location yet to be decided but under consideration for either Italy or the United States. [^37] Ibid., p. 18.
@@
One could argue that a motive behind the Italian physicists’ eagerness to collaborate with IUPAP was to mend Italy’s reputation as an Axis Power during the war. In the period following World War II, Italian physicists strongly engaged in the effort to rebuild national research infrastructures and communities, if not to build something totally new in terms of research practices and big-science-style enterprises. This attempt was based on strong cooperation among various individual and institutional actors and had in the process of internationalization a fundamental component [6]. Key figures in this (re-)construction effort, notably Edoardo Amaldi and Giovanni Polvani, collaborated to reposition Italy at the heart of the systems of international cooperation that were emerging at the time. Upon his election as president of SIF in 1947, Polvani, alongside the new SIF Council, embarked on a mission to internationalize the society’s leading journal, Il Nuovo Cimento, through innovative editorial policies and “propaganda" strategies.393939“Minutes of the Meeting of the SIF Presidential Council, October 15, 1950,” ASIF. This drive to revitalize Italian physics through international engagement prominently featured IUPAP as a pivotal ally. [^40] See Minutes of the SIF Council from the 1950s, ASIF.
@@
Italy’s prompt and engaged involvement in IUPAP since with the first post-World War II General Assembly led to the election of Amaldi to its Executive Committee during the second General Assembly in 1948. In turn, Amaldi’s role in the IUPAP Executive Committee facilitated SIF in forging a special cooperative relationship with IUPAP. While the Italian committee for IUPAP was a group appointed within the CNR Committee for Physics and Mathematics, it was SIF that pushed to promote events with financial or, in some cases, even just “moral" support from IUPAP.414141“Minutes of the SIF Presidential Council Meeting, September 24, 1952,” ASIF. This support was chiefly directed towards hosting international conferences on specific research themes in Italy, an activity intimately related to the ambition of elevating Il Nuovo Cimento to the status of an international journal and enhancing its stature within the global physics publication sphere. This is the context that led SIF to co-organize together with the Commission on Thermodynamics and Statistical Mechanics the meeting in Florence. This conference represented one of the initial and most significant efforts by SIF and Italian physicists to reassert their presence and secure a pivotal role in the realm of international scientific cooperation in the period immediately following World War II.
@@
The suggestion to host the following conference of IUPAP Commission III (then still called Commission on Thermodynamic Notations and Quantities) in Italy was already put forward during the 1948 symposium convened by Prigogine in Brussels. Among the attendees was Roberto Piontelli, an Italian professor of electrochemistry from the University of Milan, who participated as a discussant. During the symposium, Piontelli publicly proposed that the next conference of the Commission could be held in Italy with the cooperation of SIF, whose President, Polvani, was a close colleague of him at the University of Milan. Upon his return to Italy, Piontelli presented the idea to Polvani who immediately recognized the significant potential it presented. Polvani was keen to bring the event to Italy, understanding that doing so would firmly integrate Italian physicists into the global network of IUPAP-endorsed international scientific conferences [47].
@@
Amaldi proposed hosting the conference in Florence, one of the most beautiful Italian historical cities, in May, during the “Maggio Musicale Fiorentino," an internationally renowned opera and arts festival [47]. After having secured the financial support of the CNR, which allocated funds to defray the local conference costs, including hotel expenses for international speakers, [^42] Segreteria Generale del CNR to Polvani, June 30, 1948, Box 3, folder 3, Presidenza Polvani, Correspondence 1947-1959 (hereafter Polvani-ASIF), ASIF. SIF formally nominated Florence as the conference location. The enchanting Florentine setting was certainly perceived by both Italian and international physicists as an ideal backdrop to foster one of the conference’s primary objectives: the formation of an international scientific community. Initial doubts as to whether to hold the conference in Italy or the United States rapidly faded away. Florence was too attractive as a stage to establish human contacts, which, as many historians have shown, is among the main goals of scientific conferences [8, 7, 15, 55]. IUPAP President Henrik Kramers made it quite explicit in his inaugural discourse of the conference that Florence, as a symbol of Italian cultural renaissance, was a perfect place to support the renaissance of human interactions among physicists after the ruins of World War [30]. As Polvani wrote to the Rector of the University of Florence, Bruno Borghi, the conference would “be truly worthy both from a scientific point of view (the names of those who will be taking part ensure this) and from the point of view of the decorum, dignity and value that Florence and its University have unforgettably had.” [^43] Polvani to Borghi, January 27, 1949, Carteggio e Atti dell’amministrazione centrale, Filza anno 1949, Fascicolo n.19/c, “Convegno Internazionale di Meccanica Statistica,” ASUFi, University of Florence (hereafter 19/c-ASUFi). Other factors, such as the logistical ease for European scientists to travel to Italy compared to crossing the Atlantic and the strategic desire to reintegrate Italy into the circle of esteemed democratic and scientifically progressive nations while hastening the disassociation from its Fascist past, may have also influenced the decision. The allure of the SIF proposal was so compelling that by November 1948, the IUPAP Executive Committee not only embraced the invitation but Fleury also suggested scheduling its annual meeting in Florence immediately before the conference of the Commission on Thermodynamics and Statistical Mechanics, so that Executive Committee members could attend the scientific conference. [^44] Fleury to Polvani, November 27, 1948, Box 1, Folder 2, Polvani-ASIF.
@@
The Italian organizers carefully prepared for the conference, aiming to showcase the venue as exceptionally beautiful and enjoyable to foster the re-establishment of professional and personal relationships following World War II. The primary venue was the National Museum of the History of Science, now known as Museo Galileo. [^45] The conference was held in the loggia of the Castellani Palace, now the library of Museo Galileo. At that time such space was available to the homeland history delegation and it was sometimes used by the National Museum of the History of Sciences for organizing events. Bruno Borghi secured the museum’s grand lodge to properly “accomodate high-profile personalities from the scientific world, including Nobel Prize winners.” [^46] Borghi to President of the Provincial Deputation of National History, May 4, 1949, Folder “Convegno internazionale di meccanica statistica, 1949,” Fondo Corsini, Materiale minore Corsini I, Archive of the institute and Museum of the History of Science, now Museo Galileo (hereafter AIMSS) [^47] While Borghi used the plurals, the Florence conference would be attended by one Nobel Laureate at the time (Wolfgang Pauli). Three of the speakers would receive the Nobel Prize after the conference, one in physics (Max Born) and two in chemistry (Lars Onsager and Ilya Prigogine). This effort was part of a broader initiative to elevate the venue’s prestige and highlight Italy’s significant contributions to modern science’s evolution. Among the notable displays was one of the most celebrated portraits of Galileo by Justus Sustermans, on loan from the Galleria Palatina for the conference’s duration. [^48] Soprintendenza alle Gallerie per le Province di Firenze, Arezzo e Pistoia, Verbale della provvisoria consegna del ritratto di Galileo Galilei al Museo delle Science di Firenze, May 17, 1949, AIMSS. The museum’s director and staff dedicated themselves fully to the event’s success. [^49] Polvani to Corsini, May 27, 1949, AIMSS. Conference attendees enjoyed complimentary access to Florence’s major galleries and museums, such as the Galleria degli Uffizi and Galleria Pitti. They were also treated to a classical music concert, an artistic tour, and, unsurprisingly, high-quality local cuisine, both at the venue and in the charming suburb of Fiesole. [^50] Polvani to Borghi, April 15, 1949, 19/c-ASUFi; and Borghi to Azienda Autonoma del Turismo, June 5, 1949, AIMSS; and Invitation to the Gala lunch at the Aurora restaurant in Fiesole, May 17, 1949, 19/c-ASUFi. The dinner, in particular, was highlighted by Polvani for its role in offering Italian participants tangible benefits and a unique opportunity to share a communal space with international guests. [^51] Polvani to Borghi, April 15, 1949, 19/c-ASUFi.
@@
The strategic role of internationalization in the reconstruction of Italy’s national physics community is illustrated through three key initiatives undertaken by SIF in conjunction with the Florence conference. First, this international gathering served as a catalyst for SIF to secure funding from the University of Florence and other local entities, aimed at providing two substantial grants. These grants were designed to support extended research stays, lasting no less than 90 days, for two early-career SIF members at prestigious institutions abroad. Dubbed “Borse Città di Firenze," these scholarships were established by Bruno Borghi, the Rector of the University of Florence. The allocated funds for each grant were notable for the era, amounting to 500,000 Italian Lire apiece (approximately equivalent to 10,000 Euros in today’s currency) [2]. [^52] For the revaluation of currencies, we used the calculation tool in https://rivaluta.istat.it/. SIF’s strategy of leveraging conferences as opportunities to secure local funding for research grants had previously proven effective during national congresses. However, the international prominence of the Florence conference enabled SIF to significantly enhance the success of this initiative, securing a remarkable increase in funds. This underscores the pivotal role of the Florence conference in advancing Italian physics, particularly in facilitating international training. The conference’s international stature and the distinction of its attendees, including the 1945 Nobel Laureate Wolfgang Pauli (Fig. 3), were instrumental in highlighting the event’s significance and in drawing substantial support for the field within Italy.
@@
Initially, Polvani conceived the Florence research grants as a strategic initiative to bolster theoretical physics studies broadly. He proposed not to limit the research topics sponsored by the grant to statistical mechanics, envisioning that this choice would have open more opportunities to young Italian physicists. [^53] Polvani to Borghi, January 27, 1949, 19/c-ASUFi. While at least one SIF board member supported this view, [^54] Rostagni to Polvani, April 30,1949, Box 3, Folder, Polvani-ASIF. further discussion within the SIF Council led to reconsider this initial goal. In its final wording, the call declared that the grants were designed to enhance the recipients’ expertise in theoretical physics, with a particular emphasis on statistical mechanics. Eventually, the funds were allocated to support research trips explicitly tied to statistical mechanics, showing the commitment to advance Italian scholarship in this specialized area. [^55] One of the two recipients was Giorgio Careri for supporting his studies in statistical mechanics on a new method of computing the equation of state with Joseph E. Mayer at the Institute for Nuclear Studies at the University of Chicago, see Giorgio Careri to Presidenza della Società Italiana di Fisica, July 26, 1949, Box 2, Folder 2; Careri to Polvani, July 15, 1950; and Joseph E. Mayer to Italian Physical Society, Box 3, Folder 2, Polvani-ASIF. Careri’s research with Mayer was published in [38].
@@
Figure 3: Return journey from the Florence Conference, probably at the Milano Centrale railway station. From left to right: Joaquin M. Luttinger, Franca Pauli, Max R. Schafroth and Wolfgang Pauli. Courtesy CERN, Geneva, PAULI-ARCHIVE-PHO-073.

### v3

@@ IX Radiometeorology Mixed URSI, IUGG, IUPAP
3232footnotetext: Data taken from “Report 5th General Assembly, September 1947," p. 5, Series B2aa, Vol. 1, IuG. In 1947, the explicit mention of the type General had not yet been introduced, and the types Restricted and Grand were later renamed Specialized and Affiliated, respectively.
These newly established commissions fell into three categories. The first, known as “mixed commissions," were inter-union commissions created by ICSU to tackle urgent issues by linking multiple scientific unions, including IUPAP, which ratified their creation and the participation of its members. The second category, termed “grand commissions," focused on distinct sub-fields of physics. The first of this type was the International Commission for Optics (ICO), founded during the 1947 General Assembly. [^33] For the foundation of the ICO, see [37]. Its secretary was Pierre Fleury, a prominent French optical physicist who was also appointed secretary general of IUPAP at the same meeting. The establishment of the ICO set the precedent for the structure of grand commissions: they were independent international scientific entities that became integrated within IUPAP to pursue their objectives under its auspices. Grand commissions would immediately be renamed “affiliated commissions," which exist to this day, but had a limited development at the time, with the second affiliated commission established only in 1974.343434For the establishment of the second IUPAP affiliated commission, see [47]. The third category, “restricted commissions," included the Commission on Thermochemistry (Commission III), the Commission on Cosmic Rays, and the Commission on Radioactive Units. These commissions were envisioned to pursue a relatively “limited program" compared to the broader scope of mixed and, especially, grand commissions, which aimed to forge specific physics sub-fields in the international arena. Some restricted commissions were anticipated to evolve into either mixed or grand commissions. For instance, the Commission on Radioactive Units soon transitioned into the Joint Commission on Radioactivity, a collaborative effort between IUPAP and IUPAC. [^35] For a historical analysis of the Joint Commission on Radioactivity, see [25].
@@
In January 1948, under the auspices of IUPAP a first scientific meeting was organized by Prigogine at the Free University of Brussels, where he had been appointed professor of chemistry one year earlier. Entitled Symposium on Thermodynamics, the Brussels meeting saw the participation of 22 Western European chemists and physicists. They convened to explore topics in statistical thermodynamics, cryogenics, and irreversible processes.383838“IUPAP, Minutes Sixth General Assembly (1948),” p. 6, Series B2aa, Vol. 1, IuG. Some of the central topics of this symposium were connected to Prigogine’s own research agenda at the time, which related to the thermodynamics of irreversible processes. It also provided a first international opportunity to share knowledge resulting from wartime research that had previously not been properly disseminated. The deliberations and outcomes of this meeting were later disseminated in the conference proceedings, published with UNESCO’s support [66]. This event marked a foundational step in the Commission’s expanded role and scope, evolving from a body primarily focused on supporting the SUN Commission’s standardization efforts in thermodynamics to a thematic commission. This new role designated it as an IUPAP commission dedicated to fostering the development of a physics sub-discipline on an international scale, primarily through the cultivation of a dedicated scientific community. The Brussels conference, as the Commission’s inaugural activity, underscored this strategic pivot. The chosen approach to achieve international community-building was the orchestration of conferences centered on specialized research areas.
@@
This modification of the Commission’s functions was ratified at the sixth IUPAP General Assembly, convened in Amsterdam in July 1948, marking the second assembly after the end of World War II. At this meeting, the composition of the Commission was significantly broadened. Alongside Guggenheim, who served as president, and Prigogine, who took on the role of secretary, the original members, Bauer and de Boer, were joined by two scientists from the United States: physical chemist James Alexander Beattie from MIT and chemical physicist Joseph Edward Mayer from the University of Chicago.393939“IUPAP, Minutes Sixth General Assembly (1948),” p. 6, Series B2aa, Vol. 1, IuG. This inclusion of U.S. scientists represented a notable shift, as they had been absent from the first post-World War II IUPAP General Assembly and so became active in IUPAP’s commissions only from the Amsterdam General Assembly onward.
@@
The decisions regarding the Commission’s revamped function proved to be even more pivotal. Following Prigogine’s presentation of the Commission’s activities, the General Assembly endorsed a significant expansion of its responsibilities and introduced a new designation: Commission on Thermodynamics and Statistical Mechanics. [^40] Ibid., p. 13. The narrower focus on defining magnitudes and notations within thermodynamics was agreed to continue under the collaboration with the SUN Commission. This effectively reinstated the SUN Commission’s role in standardizing thermodynamic measures, as evidenced by its report at the 1948 General Assembly. Furthermore, there was an ambitious plan to host an international symposium addressing “current problems in statistical mechanics.” [^41] Ibid., p. 6. This event was tentatively scheduled for 1949, with the location yet to be decided but under consideration for either Italy or the United States. [^42] Ibid., p. 18.
@@
One could argue that one of the motivations behind Italian physicists’ eagerness to collaborate with IUPAP was to improve Italy’s reputation, tarnished by its role as an Axis Power during World War II. In the period following the war, the Italian scientific community faced major challenges such as wartime destruction and scarce resources. Additionally, the significant loss of scientists, which had begun in the 1930s, further compounded these difficulties. Many researchers had left the country, primarily for the United States, while potential young researchers often opted for different career paths. The destabilization of the scientific community began with the university professors’ oath of allegiance to fascism in the 1930s [72], and was intensified by the 1938 Fascist racial laws. At the end of the war, Italian physicists were deeply committed to overcoming these difficulties as they worked to rebuild national research infrastructures and communities. Their efforts aimed not only at restoration but also at pioneering new research practices and initiating large-scale scientific enterprises. This attempt was based on strong cooperation among various individual and institutional actors and had in the process of internationalization a fundamental component [4].
Key figures in this (re-)construction effort were Edoardo Amaldi and Giovanni Polvani. A former student and collaborator of Enrico Fermi, Amaldi took on the task of organizing the discipline after the 1938 Racial Laws dismantled the Italian physics community, leading to the departure of many leading physicists, including Fermi, Bruno Rossi, and others. After World War II, Amaldi reaffirmed his commitment by declining a position at the University of Chicago to remain in Italy. He played a key role in rebuilding Italian physics and laid the groundwork for greater European cooperation in physics [23]. [^44] Amaldi would be one of the main proponents of the European Organization of Nuclear Research (CERN) established in 1953, and of the European institutional cooperation on space research in the 1960s, see [35, 45]. Older than Amaldi, Polvani had been the director of the Institute of Physics at the University of Milan since 1929. During the 1930s, he transformed the institute into one of the centers for the development of modern physics in Italy. Like Amaldi, Polvani faced challenges in maintaining high-level research continuity at the institute after the 1938 diaspora and with limited resources. By focusing on cosmic-ray research, Polvani provided an opportunity to train a new generation of physicists, preparing them for the challenges of the post-World War II period [30].
@@
They collaborated to reposition Italy at the heart of the systems of international cooperation that were emerging at the time. Upon his election as president of SIF in 1947, Polvani, alongside the new SIF Council, embarked on a mission to internationalize the society’s leading journal, Il Nuovo Cimento, through innovative editorial policies and “propaganda" strategies.454545“Minutes of the Meeting of the SIF Presidential Council, October 15, 1950,” ASIF. This drive to revitalize Italian physics through international engagement prominently featured IUPAP as a pivotal ally. [^46] See minutes of various SIF Council’s meetings in the late 1940s and 1950s, ASIF. Italy’s prompt and engaged involvement in IUPAP since the first post-World War II General Assembly led to the election of Amaldi to its Executive Committee during the second General Assembly in 1948. In turn, Amaldi’s role in the IUPAP Executive Committee facilitated SIF in forging a special cooperative relationship with IUPAP.
@@
While the Italian committee for IUPAP was a group appointed within the CNR Committee for Physics and Mathematics, it was SIF that pushed to promote events with financial or, in some cases, even just “moral" support from IUPAP.474747“Minutes of the SIF Presidential Council Meeting, September 24, 1952,” ASIF. This support was chiefly directed towards hosting international conferences on specific research themes in Italy, an activity intimately related to the ambition of elevating Il Nuovo Cimento to the status of an international journal and enhancing its stature within the global physics publication sphere. This is the context that led SIF to co-organize together with the Commission on Thermodynamics and Statistical Mechanics the meeting in Florence. This conference represented one of the initial and most significant efforts by SIF and Italian physicists to reassert their presence and secure a pivotal role in the realm of international scientific cooperation in the period immediately following World War II.
@@
The suggestion to host the following conference of IUPAP Commission III (then still called Commission on Thermodynamic Notations and Quantities) in Italy was already put forward during the 1948 symposium convened by Prigogine in Brussels. Among the attendees was Roberto Piontelli, an Italian professor of electrochemistry from the University of Milan, who participated as a discussant. During the symposium, Piontelli publicly proposed that the next conference of the Commission could be held in Italy in cooperation with SIF, whose president, Polvani, was a close colleague of him at the University of Milan. Upon his return to Italy, Piontelli presented the idea to Polvani who immediately recognized its significant potential. Polvani was eager to bring the event to Italy, understanding that it would firmly integrate Italian physicists into the network of IUPAP-endorsed international scientific conferences [65].
Amaldi proposed hosting the conference in Florence, one of the most beautiful Italian historical cities, in May, during the “Maggio Musicale Fiorentino," an internationally renowned opera and arts festival [65]. After having secured the financial support of the CNR, which allocated funds to defray the local conference costs, including hotel expenses for international speakers, [^48] Segreteria Generale del CNR to Polvani, June 30, 1948, Box 3, Folder 3, Presidenza Polvani, Correspondence 1947-1959 (hereafter Polvani-ASIF), ASIF. SIF formally nominated Florence as the conference location.
@@
The enchanting Florentine setting was certainly perceived by both Italian and international physicists as an ideal backdrop to foster one of the conference’s primary objectives: the formation of an international scientific community. Initial doubts as to whether to hold the conference in Italy or the United States rapidly faded away. Florence was too attractive as a stage to establish human contacts, which, as many historians have shown, is among the main goals of scientific conferences [6, 5, 20, 76]. IUPAP President Henrik Kramers made it quite explicit in his inaugural discourse of the conference that Florence, as a symbol of Italian cultural renaissance, was a perfect place to support the renaissance of human interactions among physicists after the ruins of World War [44]. As Polvani wrote to the Rector of the University of Florence, Bruno Borghi, the conference would “be truly worthy both from a scientific point of view (the names of those who will be taking part ensure this) and from the point of view of the decorum, dignity and value that Florence and its University have unforgettably had.” [^49] Polvani to Borghi, January 27, 1949, Carteggio e Atti dell’amministrazione centrale, Filza anno 1949, Fascicolo n.19/c, “Convegno Internazionale di Meccanica Statistica,” ASUFi, University of Florence (hereafter 19/c-ASUFi). Other factors, such as the logistical ease for European scientists to travel to Italy compared to crossing the Atlantic and the strategic desire to reintegrate Italy into the circle of esteemed democratic and scientifically progressive nations while hastening the disassociation from its Fascist past, may have also influenced the decision. The allure of the SIF proposal was so compelling that by November 1948, the IUPAP Executive Committee not only embraced the invitation but Fleury also suggested scheduling its annual meeting in Florence immediately before the conference of the Commission on Thermodynamics and Statistical Mechanics, so that Executive Committee members could attend the scientific conference. [^50] Fleury to Polvani, November 27, 1948, Box 1, Folder 2, Polvani-ASIF.
@@
The Italian organizers carefully prepared for the conference, aiming to showcase the venue as exceptionally beautiful and enjoyable to foster the re-establishment of professional and personal relationships following World War II. The primary venue was the National Museum of the History of Science, now known as Museo Galileo. [^51] The conference was held in the loggia of the Castellani Palace, now the library of Museo Galileo. At that time such space was available to the homeland history delegation and it was sometimes used by the National Museum of the History of Sciences for organizing events. Bruno Borghi secured the museum’s grand lodge to properly “accomodate high-profile personalities from the scientific world, including Nobel Prize winners.” [^52] Borghi to President of the Provincial Deputation of National History, May 4, 1949, Folder “Convegno internazionale di meccanica statistica, 1949,” Fondo Corsini, Materiale minore Corsini I, Archive of the Institute and Museum of the History of Science, now Museo Galileo (hereafter AIMSS). [^53] While Borghi used the plurals, the Florence conference would be attended by one Nobel Laureate at the time (Wolfgang Pauli). Three of the speakers would receive the Nobel Prize after the conference, one in physics (Max Born) and two in chemistry (Lars Onsager and Ilya Prigogine). This effort was part of a broader initiative to elevate the venue’s prestige and highlight Italy’s significant contributions to modern science’s evolution. Among the notable displays was one of the most celebrated portraits of Galileo by Justus Sustermans, on loan from the Galleria Palatina for the conference’s duration. [^54] Soprintendenza alle Gallerie per le Province di Firenze, Arezzo e Pistoia, Verbale della provvisoria consegna del ritratto di Galileo Galilei al Museo delle Scienze di Firenze, May 17, 1949, AIMSS. The museum’s director and staff dedicated themselves fully to the event’s success. [^55] Polvani to Corsini, May 27, 1949, AIMSS. Conference attendees enjoyed complimentary access to Florence’s major galleries and museums, such as the Galleria degli Uffizi and Galleria Pitti. They were also treated to a classical music concert, an artistic tour, and, unsurprisingly, high-quality local cuisine, both at the venue and in the charming suburb of Fiesole. [^56] Polvani to Borghi, April 15, 1949, 19/c-ASUFi; Borghi to Azienda Autonoma del Turismo, June 5, 1949, AIMSS; and Invitation to the Gala lunch at the Aurora restaurant in Fiesole, May 17, 1949, 19/c-ASUFi. The dinner, in particular, was highlighted by Polvani for its role in offering Italian participants tangible benefits and a unique opportunity to share a communal space with international guests. [^57] Polvani to Borghi, April 15, 1949, 19/c-ASUFi.
@@
The strategic role of internationalization in the reconstruction of Italy’s national physics community is illustrated through three key initiatives undertaken by SIF in conjunction with the Florence conference. First, this international gathering served as a catalyst for SIF to secure funding from the University of Florence and other local entities, aimed at providing two substantial grants. These grants were designed to support extended research stays, lasting no less than 90 days, for two early-career SIF members at prestigious institutions abroad. Dubbed “Borse Città di Firenze," these scholarships were established by the Rector of the University of Florence. The allocated funds for each grant were notable for the era, amounting to 500,000 Italian Lire apiece (approximately equivalent to 10,000 Euros in today’s currency) [3]. [^58] For the revaluation of currencies, we used the calculation tool in https://rivaluta.istat.it/. SIF’s strategy of leveraging conferences as opportunities to secure local funding for research grants had previously proven effective during national congresses. However, the international prominence of the Florence conference enabled SIF to significantly enhance the success of this initiative, securing a remarkable increase in funds. This signals the pivotal role of the Florence conference in advancing Italian physics, particularly in facilitating international training. The conference’s international stature and the distinction of its attendees, including the 1945 Nobel Laureate Wolfgang Pauli (Fig. 3), were instrumental in highlighting the event’s significance and in drawing substantial support for the field within Italy.
@@
Initially, Polvani conceived the Florence research grants as a strategic initiative to bolster theoretical physics studies broadly. He proposed not to limit the research topics sponsored by the grant to statistical mechanics, envisioning that this choice would have open more opportunities to young Italian physicists. [^59] Polvani to Borghi, January 27, 1949, 19/c-ASUFi. While at least one SIF board member supported this view, [^60] Rostagni to Polvani, April 30,1949, Box 3, Folder, Polvani-ASIF. further discussion within the SIF Council led to reconsider this initial goal. In its final wording, the call declared that the grants were designed to enhance the recipients’ expertise in theoretical physics, with a particular emphasis on statistical mechanics. Eventually, the funds were allocated especially to support research trips tied to statistical mechanics, showing the commitment to advance Italian scholarship in this specialized area. One of the two recipients was Giorgio Careri, who was supported his studies on the presence of a statistical order in the condensed phase, conducted in collaboration with Mayer at the Institute for Nuclear Studies at the University of Chicago. [^61] Giorgio Careri to Presidenza della Società Italiana di Fisica, July 26, 1949, Box 2, Folder 2; Careri to Polvani, July 15, 1950; and Joseph E. Mayer to Italian Physical Society, Box 3, Folder 2, Polvani-ASIF. Careri’s research stay in Chicago was further supported by a grant of the Fulbright Program obtained with the support of Amaldi [10, pp. 173-84]. Careri’s research with Mayer was published in [55]. A few years after his return, Careri established schools focused on the properties of liquid helium and low temperature physics, topics that were novel in the Italian physics landscape [11, pp. 113-43] [see also 9]. This initiative significantly enhanced Italian research in areas related to statistical mechanics and condensed matter physics. In later recollections, Careri himself emphasized the pivotal role of the Florence conference in the opening up of these new research avenues in Italy [16, 17]. [^62] The second recipient was Giampiero Puppi; see “Minutes of the SIF Presidential Council Meeting, September 3, 1949,” ASIF. However, it in unclear whether Puppi actually made use of the funds and, eventually, his research interests remained strictly related to theoretical particle physics.
@@
Figure 3: Return journey from the Florence Conference, probably at the central railway station of Milan. From left to right: Joaquin M. Luttinger, Franca Pauli, Max R. Schafroth and Wolfgang Pauli. Courtesy CERN, Geneva, PAULI-ARCHIVE-PHO-073.